OpenStudy (anonymous):

Daniela has $35.50 in dimes and quarters. She has 5 more dimes than quarters. How many quarters does she have?

OpenStudy (lgbasallote):

let d= dimes q = quarters 5 more dimes than quarters that means d = 5 + q total money is 35.50 a dime is 10 cents a quarter is 25 cents $35.50 is 3550 cents 10d + 25q = 3550 now remember d = 5 + q <--sub that 10(5+q) + 25q = 3550 are you good now?

OpenStudy (anonymous):

a lot?

OpenStudy (anonymous):

yea thanks

OpenStudy (lgbasallote):

great!! ^_^